OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)?

OpenStudy (rosedewittbukater):

Slope is rise/run, so do y2-y1/x2-x1 -2-4/2-(-2) = -6/2+2 = -6/4 = -3/2 The slope is \[-\frac{ 3 }{ 2 }\]